Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group November 2, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 2, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/02/98 (Monday)

Segment 1 (Emidio to ANF): This report has not been received.

Segment 2 (ANF to Sylmar) Report for 11/02/98: 2A (Copco Rd to Templin): The Liebre testing cutout, EMP 42.27, was tied-in. The Forest Inn cutout has also been tied-in and backfilled. 2C (Templin Hwy to Castaic): Construction began today on a block wall on the Whitaker Reduction Station access road. The testing cutout south of Whitaker Station, EMP 51.50, was backfilled. F/O cable was pulled north of the I-5 bore, EMP 53.45. Slope stabilization and revegetation occurred near EMP 53.38 and 54.55, and south of EMP 54.74. Welding moved north on the Edison relocation to EMP 55.88. Backfill reached 56.06. 2B (Castaic to Sylmar): At the intersection of The Old Road and Hwy 126, F/O cable was spliced. At Pico Cyn. on The Old Road, an assist pit was dug to aid in pulling the F/O cable. At Calgrove on The Old Road, the north bore pit was excavated. Just north of Sunshine Cyn., 500 feet of coating, 330 feet of base rock application, and 220 feet of F/O conduit-laying was accomplished. At Sunshine Cyn., the riser was welded and x-rayed. At Balboa Blvd, slurry was poured under the bridge around the vault, and trench for an electric line was excavated across San Fernando Rd.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 2 November 1998: 3A: The coating crew continued their work at the Verdugo Wash Bridge today. Welding, x-ray, and coating occurred at both the San Fernando and Hollywood Way Bridges. A ditch crew spent the day at the Van Nuys Boulevard bore site. The crew worked on fixing a cave in on the ditch, as well as cleaning dirt from the pipe. NIGHT work consisted of a short time for concrete cutting at Hollywood Way. 3B: Backfill progressed on their work on the north end of the Downey Bridge tie in. A crew worked from the Los Angeles River bottom to hang razor wire on the Downey Bridge. 3C: Excavation began on the bore pit for Hubbard Avenue. Clean up occurred along the right of way between Hubbard and Roxford Streets. Backfill worked on closing up the ditch near Grapevine Creek. Ditch crews progressed slowly along their narrow right of way, north of Sepulveda Street. Concrete was pumped into the frames for the California Aqueduct crossing.

Segment 4B (El Segundo to Lynwood) 11-2-98: Backfilling activities continued at 108 Th St. ESEMP 1.16, Santa Ana ESEMP 0.04, Alameda and Fernwood AV ESEMP 0.0. Coating and slurry backfilling took place at Western ESEMP 5.56. Construction activities continued at El Segundo station ESEMP 11.00. This crew was sending pigs through the casing that goes under the Sepulveda BLVD. A small crew from Hood was excavating for a tie-in at Sepulveda ROW near the El Segundo station ESEMP 10.89. Construction activities continued at Lynwood station EMP 0.00.

Segment 5 (Lynwood to Long Beach) Report for 2 November 1998: Directional drilling crew at Highway 91 had the lead cable break and had to pull back pipe to locate the end of the cable. The cable was repaired and drilling resumed. The crew at Wilmington finished the excavation of the extra-depth bore pit and are waiting for OSHA inspection before proceeding.

Inspection Summary

In the past 24 hours, 4 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page